OpenStudy (anonymous):

Find the least positive integer n so that both n and n + 1 have prime factorizations with exactly four (not necessarily distinct) prime factors.

OpenStudy (kinggeorge):

I'm not sure how to solve this mathematically. The only way I've come up with so far, is to just try different numbers starting from 16 until we hit one where this works.

OpenStudy (anonymous):

Do you have one that works??

OpenStudy (kinggeorge):

\[3^3\cdot5=135\]\[3^3\cdot5+1=136=2^3\cdot17\]Is the smallest I could find.

OpenStudy (kinggeorge):

I found this by trying numbers with 4 prime factors starting from the least and increasing by the smallest amount possible each time. This was the first I came across.

OpenStudy (anonymous):

looks great, THANKS!!!